OpenStudy (anonymous):

Solve the given differential equation by using the substitution \(u=y'\) \[y''+(y')^2+1=0, u\frac{du}{dy}=y''\] \[udu+u^2dy+dy=0\] \[udu=-u^2dy-dy=-dy(u^2+1)\] \[\int \frac{u}{u^2+1}du=-\int dy\] \[v=u^2+1,\frac{dv}{2}=udu\] \[\frac{1}{2}\int \frac{dv}{v}=-\int dy\] \[\ln(u^2+1)=-2(y+\ln(c_1))\] \[u^2=c_1^{-2}e^{-2y}-1\] \[u=\pm\sqrt{c_1^{-2}e^{-2y}-1}=\frac{dy}{dx}\] I'm not sure where I can go from here.

ganeshie8 (ganeshie8):

there is a mistake in the fist line

ganeshie8 (ganeshie8):

\(\large y' = u \implies y'' = u'\) so the equation must become become : \(\large u' + u^2 = -1\)

ganeshie8 (ganeshie8):

\(\large \dfrac{du}{dx}= -(1+u^2)\) \(\large \int \dfrac{du}{1+u^2} = \int - dx\)

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

It's looking like that will get the answer in the book. I did the substitution based on the professor mentioning using that as the chain rule, but there is no y term.

ganeshie8 (ganeshie8):

yes this substitution works only when there is no y term

ganeshie8 (ganeshie8):

DE needs to be of form f(y'', y', x) = 0

OpenStudy (anonymous):

\[y''+(y')^2+1=0, u'=y''\] \[u'+u^2+1=0\] \[\int \frac{du}{u^2+1}=-\int dx\] \[\frac{1}{2}\int \frac{dv}{v}=-\int dx\] \[\arctan(u)=c_1-x\] \[u=\tan(c_1-x)\] \[dy=\tan(c_1-x)\ \[u=\cos(c_1-x_,du=\sin(c_1-x)dx\] \[dy=\frac{du}{u}\] \[y=\ln|cos(c_1-x)|+c_2\] And that is correct.

ganeshie8 (ganeshie8):

looks good to me :)